# Superposition Yields Robust Neural Scaling

### v2

@@ Abstract
The success of today’s large language models (LLMs) depends on the observation that larger models perform better. However, the origin of this neural scaling law — the finding that loss decreases as a power law with model size — remains unclear. Starting from two empirical principles — that LLMs represent more things than the model dimensions (widths) they have (i.e., representations are superposed), and that words or concepts in language occur with varying frequencies — we constructed a toy model to study the loss scaling with model size. We found that when superposition is weak, meaning only the most frequent features are represented without interference, the scaling of loss with model size depends on the underlying feature frequency; if feature frequencies follow a power law, so does the loss. In contrast, under strong superposition, where all features are represented but overlap with each other, the loss becomes inversely proportional to the model dimension across a wide range of feature frequency distributions. This robust scaling behavior is explained geometrically: when many more vectors are packed into a lower dimensional space, the interference (squared overlaps) between vectors scales inversely with that dimension. We then analyzed four families of open-sourced LLMs and found that they exhibit strong superposition and quantitatively match the predictions of our toy model. The Chinchilla scaling law turned out to also agree with our results. We conclude that representation superposition is an important mechanism underlying the observed neural scaling laws. We anticipate that these insights will inspire new training strategies and model architectures to achieve better performance with less computation and fewer parameters.
@@ 1 Introduction
Figure 1: Superposition leads to robust and fast power-law loss decay with model size. (a) Illustration of no superposition where a three-dimensional space can at most represent three features without any interference (overlap). (b) Toy model results in the regime of weak superposition, where we set data dimension $n=10240$ (number of features). The toy model will be introduced; more details are in Appendix D.1. (c) Illustration of superposition: there are more features than the dimension of the space. (d) The same toy models in the strong superposition regime show lower losses, which are on power laws with model dimension and have exponents close to $1$ (color coding same as panel b). The gray points are from actual LLMs, which seem to have a similar power-law relation with exponent $\sim 1$.
@@
The scaling of loss with model size plays a central role in both the practical design and the theoretical understanding of large-scale machine learning systems, yet its origin remains inconclusive [3, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Various explanations have been proposed, drawing from statistical learning theory and empirical phenomenological models, including improved function or manifold approximation in larger models [14, 15], and enhanced representation or skill learning in larger models [18, 19, 20, 21]. In the limit of infinite data, many of these explanations predict a power-law decay of loss with model size, provided the underlying data distribution also follows a power law. However, the predicted scaling exponents depend sensitively on the properties of the data distribution. Moreover, the connection between these mechanistic explanations and the behavior of actual LLMs needs further exploration.
@@
When considering LLMs specifically, it becomes clear that representation learning can be a limiting factor, which is closely related to a phenomenon called superposition [25, 26], yet this aspect has not been thoroughly studied. LLMs must learn embedding vectors for tokens, process these representations through transformer layers to predict the next token, and use a final projection (the language model head) to generate the output. Conceptually, fitting functions or manifolds and learning skills or grammars are primarily tasks of the transformer layers, while representation learning is more directly tied to the embedding matrix and the language model head. To represent more than fifty thousand tokens — or even more abstract concepts — within a hidden space of at most a few thousand dimensions, the quality of representations is inevitably constrained, contributing to the final loss. Although models can represent more features than their dimensionality would suggest through a mechanism known as superposition [26], prior work on neural scaling laws has implicitly assumed weak superposition [15, 16, 17, 18, 19], which may be less relevant to the regime where LLMs operate. To address this gap, we systematically study here how representation superposition and data structure together influence the scaling of loss with model size.
@@
We adopt a toy model construction similar to [26] to study how superposition affects neural scaling laws. In the toy model, representations are learned by recovering data, each composed of multiple latent features. These features in data have different frequencies of occurrence, reflecting their relative importance. In the absence of superposition, only the most frequent features are perfectly represented, while the others are ignored. As illustrated in Figure 1a, the first three of six features are represented in the three-dimensional space without interference, and the remaining three are omitted. We find that in the weak superposition regime, the scaling of loss with model dimension depends sensitively on how feature frequency decays with rank: the loss follows a power law with model size only if the feature frequencies themselves follow a power law, provided that $m$ is sufficiently large (Figure 1b). By contrast, strong superposition allows many more features to be represented, albeit with overlap in the representation (Figure 1c). In this regime, the model displays a robust behavior: loss scales inversely with model dimension, independent of the data frequency distribution (Figure 1d). Remarkably, we find that actual LLMs follow the same rule. Comparing models with the same dimension but different superposition degrees, the ones with stronger superposition also have a lower loss. Building on the basic phenomena, we provide the main takeaways: (i) In the strong superposition regime, loss due to imperfect representation scales inversely with model dimension, largely independent of feature frequency distributions; (ii) Scaling behaviors in the weak superposition regime are brittle and depend on the number of activated but unlearned features, and in the strong superposition regime are robust and can be understood via generic geometric facts; (iii) LLMs exhibit strong superposition and agree quantitatively with our toy model predictions.
@@
The rest of the paper will elaborate on the takeaways. In Section 2, we introduce the toy model, describe the data sampling procedure, and explain how we control the degree of superposition. Section 3 presents the detailed results and quantitative analysis of how superposition and data distribution influence loss scaling. In Section 4, we compare our findings to related works. Finally, Section 5 summarizes our conclusions and discusses limitations and future directions. Codes are available at https://github.com/liuyz0/SuperpositionScaling.
@@ 2 Methods
Figure 2: Toy model of superposition. (a) Architecture and loss of the toy model. (b and c) Rows of the weight matrix $W$, denoted by $W_{i}$, are representations for features. (b) No superposition represented the most frequent features, i.e., the first three, without interference ($WW^{T}$ has zero off-diagonal elements). In the frequency-rank plot, height means frequency, and color means $\|W_{i}\|_{2}$. (c) With superposition, more features are represented, while the vectors $W_{i}$ overlap.
To gain understanding of the relationship between superposition and data structure, we need a toy model on learning representation simple enough yet not simpler — two key principles need to be reflected, (i) there are more features to represent than the dimension of the model, and (ii) features occur in data with different frequencies. Later, we will discuss how the loss due to representation studied here may affect the overall final loss in LLMs.
We adopt the toy model of superposition from Anthropic [26] with minor modifications (Figure 2a). Input data $x\in\mathbb{R}^{n}$ is a vector with data dimension $n$ being the number of atomic (or irreducible) features. Each element $x_{i}$ in $x$ is interpreted as the activation of this data point at feature $i$, which is sampled as
@@
Here, $u_{i}$ sampled from a Bernoulli distribution controls whether the feature $i$ is activated, and $v_{i}$ sampled from a uniform distribution controls the activation strength once feature $i$ is activated. The frequency of feature $i$ to appear in the data is $p_{i}$. Without loss of generality, we make the indices of features the same as their frequency or importance rank. The data structure is then about how $p_{i}$ decreases with rank $i$. The expected number of activations in one data will be referred to as activation density: $E=\sum_{i=1}^{n}p_{i}$. The model learns hidden representations by recovering the data, which cannot be done perfectly because the model dimension $m$ is much smaller than the number of possible features in the data $n$. The trainable parameters are a weight matrix $W\in\mathbb{R}^{n\times m}$ and a bias vector $b\in\mathbb{R}^{n}$. The weight matrix embeds data $x$ into a hidden space with dimension $m$, $h=W^{T}x$, with $m\ll n$. In practice, we fix $n$ as a large number and change the model dimension $m$. We use $W$ to read out the embedding, where $y=\mathrm{ReLU}(Wh+b)$. The loss is defined as the difference between the recovered $y$ and the original $x$, $L=\langle\|y-x\|_{2}^{2}\rangle_{x}$, where $\langle\cdot\rangle_{x}$ means average over $x$ distribution.
@@
We can now formally introduce superposition. Note that $W_{i}$ is the representation of feature $i$ in the hidden space, where we use $W_{i}$ to denote the $i$th row of the $W$ matrix, in our toy model, feature $i$ is well represented when the norm $\|W_{i}\|_{2}$ is close to $1$. No superposition ideally means the first $m$ rows of $W$ form an orthogonal basis (i.e., the first $m$ most important features represented perfectly) and the rest of the rows are zero (i.e., rest of the features ignored or lost), as illustrated in Figure 2b. Superposition means that there are more than $m$ rows in $W$ with norms close to $1$ (Figure 2c). Conceptually, superposition is preferred when features are more equal in importance ($p_{i}$ are more uniform) and when features are more sparse ($E/n$ is smaller) [26].
Figure 3: Weight decay can robustly tune the degree of superposition. (a) Given data dimension $n=1000$ and feature frequency $p_{i}\propto 1/i^{\alpha}$ with $\alpha=1$, we tested model dimension $m$ from $10$ to $100$ and weight decay from $-1$ to $1$. The frequency-rank plot shows that important features tend to be represented (color means $\|W_{i}\|_{2}$ and colorbar same as Figure 2b). The row norm distribution is bimodal, making $\phi_{1/2}$ a good measure for the fraction represented. Smaller weight decay leads to larger $\phi_{1/2}$ or stronger superposition. (b) After changing $\alpha$, the ability for small weight decay to reach strong superposition is not changed. We show the result at $m=39$ here, more in Appendix D.3.
We found that it is possible to control superposition independently of data properties by modifying the optimizer with a weight decay (or growth) term:
@@
where $\eta_{t}$ is the learning rate and $W{i,t}$ is the $i$th row of the weight matrix at step $t$ (vector operations are element-wise). For weight decay $\gamma<0$, the update corresponds to gradient descent on $(|W_{i,t}|_{2}-1)^{2}$, encouraging unit-norm rows. We implement this in a modified AdamW [27] optimizer with a warm-up and cosine decay learning rate schedule. Training runs for 20k–80k steps with batch sizes of 2048–8192, ensuring models reach saturation. At each step, we sample new data using probabilities $p_{i}$. We observe that important features are represented as expected, and row norms of $W$ become bimodal, clustering near 0 and 1 (Figure 3a). This allows us to define the fraction of represented features as
@@
namely, the fraction of rows with norm larger than $1/2$. We found that weight decay can tune superposition smoothly, with small weight decay $\gamma$ giving strong superposition, i.e., $\phi_{1/2}\approx 1\gg m/n$, and large weight decay corresponding to weak superposition, i.e., $\phi_{1/2}\sim m/n$ (Figure 3a). The ability of weight decay to tune superposition is robust to a variety of feature frequency distributions (Figure 3b), where $p_{i}\propto 1/i^{\alpha}$ and we control the decay speed of $p_{i}$ via the data exponent $\alpha$. Thus, weight decay provides a practical knob for controlling superposition in our study.
Our toy model differs from LLMs in architecture, data, and loss. Since we focus on representations rather than next-token prediction, we omit transformer layers. Conceptually, LLMs map a document to a token, with inputs and outputs in different spaces, while the toy model operates within a single shared space. Despite this, the toy model captures key aspects of language structure through engineered sparsity and feature importance, making its data structure aligned with that of LLMs at a higher level. While LLMs use cross-entropy loss and the toy model uses squared error, we will show that this does not affect the scaling behavior of the loss. Thus, the toy model is a suitable abstraction for studying representation-limited scaling.
@@ 3 Results
Figure 4: Rich scaling phenomena arise when we change the degree of superposition and data structures. (a) In the same experiments ($n=10240$) in Figure 1, where $p_{i}\propto\exp(-i/400)$ (blue), $p_{i}\propto 1/i^{1.2}$ (yellow), and $p_{i}\propto n-i$ (green), we use $\gamma=-1$ to reach strong superposition and $\gamma=0.1$ for weak superposition. R-squared value of the fitting is used to measure how likely the fitted part is a power law (Appendix D.4). R-squared values closer to 1 mean that the data are more similar to a power law. We found that at strong superposition, power laws are robust across different underlying feature distributions. Yet, at weak superposition, only power-law feature frequencies can lead to power-law losses. (b) When changing superposition by weight decay and varying feature frequency decay by $\alpha$ given $p_{i}\propto 1/i^{\alpha}$, we found roughly three distinct behaviors. For (b), $n=1000$ and $m=10\sim 100$ (more in Appendix D.4).
In the toy model, we found rich phenomena by varying superposition and feature importance. Reviewing the experiments in Figure 1, we set weight decay $\gamma=-1$ to have strong superposition and $\gamma=0.1$ to have weak superposition. We compute $R$ squared values from linear fits in log-log plots to quantify scaling behavior, assessing how closely the loss follows a power law to model dimension. We can see that at strong superposition, the losses are close to power laws, regardless of the underlying feature frequencies, yet the loss is a power law at weak superposition if the feature frequency $p_{i}$ is a power law with rank $i$ (Figure 4a). For a systematic scan, we next set $p_{i}\propto 1/i^{\alpha}$ and can vary the data exponent $\alpha$ to change how fast $p_{i}$ decays consistently. [^1] The word or phrase frequency in natural language follows Zipf’s law, which is a power law. Unless otherwise specified, we set activation density $E$ as $1$ by default, whose value will be shown not to affect the scaling. Assuming a power-law form for the final test loss, $L\propto 1/m^{\alpha_{m}}$, we extract the model exponent $\alpha_{m}$ from the empirical fit. We fit the loss with a power law in all cases. The fitted $\alpha_{m}$ reveals how fast losses decay, even in the weak superposition regime where the power law fit is poor (Appendix D.4). Roughly, three distinct patterns emerge: (1) under weak superposition (positive $\gamma$, yellow box in Figure 4b), $\alpha_{m}$ is small, indicating slow loss decay; (2) under strong superposition and a wide range of small data exponents (red box), $\alpha_{m}$ remains robustly near 1; (3) for strong superposition with large data exponents (blue box), $\alpha_{m}$ increases with $\alpha$. By interpreting these three patterns, we aim to understand when loss follows a power law with model dimension, and what determines the exponent when it does.

### v3

@@ Abstract
The success of today’s large language models (LLMs) depends on the observation that larger models perform better. However, the origin of this neural scaling law, that loss decreases as a power law with model size, remains unclear. We propose that representation superposition, meaning that LLMs represent more features than they have dimensions, can be a key contributor to loss and cause neural scaling. Based on Anthropic’s toy model, we use weight decay to control the degree of superposition, allowing us to systematically study how loss scales with model size. When superposition is weak, the loss follows a power law only if data feature frequencies are power-law distributed. In contrast, under strong superposition, the loss generically scales inversely with model dimension across a broad class of frequency distributions, due to geometric overlaps between representation vectors. We confirmed that open-sourced LLMs operate in the strong superposition regime and have loss scaling like one over the model dimension, and that the Chinchilla scaling laws are also consistent with this behavior. Our results identify representation superposition as a central driver of neural scaling laws, providing insights into questions like when neural scaling laws can be improved and when they will break down. [^1] Code is available at https://github.com/liuyz0/SuperpositionScaling
@@ 1 Introduction
Figure 1: Superposition leads to robust and fast power-law loss decay with model size. (a) Illustration of no superposition where a three-dimensional space can at most represent three features without any interference (overlap). (b) Toy model results in the regime of weak superposition, where we set data dimension $n=10240$ (number of features). The toy model will be introduced; more details are in Appendix D.1. (c) Illustration of superposition: there are more features than the dimension of the space. (d) The same toy models in the strong superposition regime show lower losses, which are on power laws with model dimension and have exponents close to 1 (color coding same as panel b). The gray points are from actual LLMs, which have a similar power-law exponent near 1.
@@
The power-law loss with model size plays a central role in both the practical design and the theoretical understanding of large-scale machine learning systems, yet its origin remains inconclusive [3, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Various explanations have been proposed, drawing from statistical learning theory and empirical phenomenological models, including improved function or manifold approximation in larger models [14, 15], and enhanced representation or skill learning in larger models [19, 20, 21, 22]. In the limit of infinite data, many of these explanations predict a power-law decay of loss with model size, provided the underlying data distribution also follows a power law. The scaling exponents are sensitive to the properties of the data distribution. Moreover, the connection between these mechanistic explanations and the behavior of actual LLMs needs further exploration.
@@
When considering LLMs specifically, it becomes clear that representation or embedding can be a limiting factor, which is closely related to a phenomenon called superposition [26, 27], yet this aspect has not been thoroughly studied. LLMs must learn embedding vectors for tokens, process these representations through transformer layers to predict the next token, and use a final projection (the language model head) to generate the output. Conceptually, fitting functions or manifolds and learning skills or grammars are primarily tasks of the transformer layers, while representation is more directly tied to the embedding matrix and the language model head. To represent more than fifty thousand tokens — or even more abstract concepts — within a hidden space of at most a few thousand dimensions, the quality of representations is inevitably constrained by the model dimension or width, contributing to the final loss. Although models can represent more features than their dimensionality would suggest through a mechanism known as superposition [27], prior works on neural scaling laws seem to fall in the weak superposition regime implicitly [15, 16, 17, 18, 19, 20], which may be less relevant to the regime where LLMs operate. This gap leads us to study
Question: How will superposition influence the loss scaling with model dimension (width)? Varying the degree of superposition and data structure, when is the loss a power law? And if the loss is a power law, what will the exponent be?
@@
We adopt a toy model construction similar to [27] to study how superposition affects neural scaling laws. In the toy model, representations are learned by recovering data, each composed of multiple latent features. These features in data have different frequencies of occurrence, reflecting their relative importance. Weak superposition means that only the most frequent features are perfectly represented, while the others are ignored. As illustrated in Figure 1a, the first three of six features are represented in the three-dimensional space without interference, and the remaining three are omitted. We find that in the weak superposition regime, the scaling of loss with model dimension depends sensitively on how feature frequency decays with rank: the loss follows a power law with model size only if the feature frequencies themselves follow a power law, provided that $m$ is sufficiently large (Figure 1b). By contrast, strong superposition allows many more features to be represented, albeit with overlap in the representation (Figure 1c). In this regime, the model displays a robust behavior: loss scales inversely with model dimension across different data frequency distributions (Figure 1d). Remarkably, we find that actual LLMs follow a similar scaling. We summarize our contributions as
Main results/messages $\bullet$ Loss in the weak superposition regime depends on summing frequencies of ignored features, which is a power law if frequencies follow a power law. $\bullet$ In the strong superposition regime, loss arises from the interference between representations and can have robust “one over width" scaling because of the geometry. $\bullet$ LLMs exhibit strong superposition and agree quantitatively with our toy model predictions.
@@
The rest of the paper will elaborate on the takeaways. In Section 2, we introduce the toy model, describe the data sampling procedure, and explain how we control the degree of superposition. Section 3 presents the detailed results. In Section 4, we compare our findings to related works. Finally, Section 5 summarizes our conclusions and discusses limitations and future directions.
@@ 2 Methods
Figure 2: Toy model of superposition. (a) Architecture and loss of the toy model. (b and c) A row of the matrix $W$, denoted by $W_{i}$, is the representation of feature ii. (b) No superposition represented the most frequent features, i.e., the first three ($n=6$ features in $m=3$ dimensional space), without interference. In the frequency-rank plot, height means feature $i$’s frequency $p_{i}$, and color means the $i$th row vector’s norm $\|W_{i}\|_{2}$. (c) With superposition, features are all represented, while the representations $W_{i}$ overlap.
To understand the relationship between superposition and data structure, we need a toy model to represent data features simple enough yet not simpler — two key principles need to be reflected, (i) there are more features to represent than the dimension of the model, and (ii) features occur in data with different frequencies. Later, we will discuss how the loss due to representation studied here may affect the overall final loss in LLMs.
We adopt the toy model of superposition from Anthropic [27] (an autoencoder) with minor modifications (Figure 2a). Input $x\in\mathbb{R}^{n}$ is a vector with data dimension $n$ being the number of atomic (or irreducible) features. Each element $x_{i}$ in $x$ is interpreted as the activation of this sample at feature $i$, which follows
@@
Here, $u_{i}$ sampled from a Bernoulli distribution controls whether the feature $i$ is activated, and $v_{i}$ sampled from a uniform distribution controls the activation strength once feature $i$ is activated. All samples are i.i.d. The frequency of feature $i$ to appear in the data is $p_{i}$. Without loss of generality, we make the indices of features the same as their frequency or importance rank. The data structure is then about how $p_{i}$ decreases with rank $i$. The expected number of activations in one input will be referred to as activation density: $E=\sum_{i=1}^{n}p_{i}$. The model learns hidden representations by recovering the data, which cannot be done perfectly because the model dimension $m$ is much smaller than the number of possible features in the data $n$. The trainable parameters are a weight matrix $W\in\mathbb{R}^{n\times m}$ and a bias vector $b\in\mathbb{R}^{n}$. The weight matrix embeds data $x$ into a hidden space with dimension $m$, $h=W^{T}x$, with $m\ll n$. In practice, we fix $n$ as a large number and change the model dimension $m$. We use $W$ to read out the embedding, where $y=\mathrm{ReLU}(Wh+b)$. The loss is defined as the difference between the recovered $y$ and the original $x$, $L=\langle\|y-x\|_{2}^{2}\rangle_{x}$, where $\langle\cdot\rangle_{x}$ means average over $x$ distribution.
We can now formally introduce superposition. Note that $W_{i}$ is the representation of feature $i$ in the hidden space, where we use $W_{i}$ to denote the $i$th row of the $W$ matrix. We emphasize the following
Key concepts $\bullet$ Feature frequency: $p_{i}$ is the probability that feature $i$ is activated (non-zero) in a sample, which is assumed to decrease with $i$. $\bullet$ Sparsity: We say features are sparse when $E/n$ is small. $\bullet$ The feature $i$ is represented (in the hidden space) when $W_{i}$ is non-zero.
@@
No superposition ideally means the first $m$ rows of $W$ form an orthogonal basis (i.e., the first $m$ most important features represented perfectly) and the rest of the rows are zero (i.e., the rest of the features ignored or lost), as illustrated in Figure 2b. Superposition means that there are more than $m$ rows in $W$ with non-zero norms (Figure 2c).
We next summarize important facts of this toy model [27]:
Preliminaries Superposition cannot lead to lower losses in a linear model (without ReLU function) due to the large amount of interference. ReLU and negative biases can cancel off interference, realizing error correction. With this non-linearity, superposition can be preferred when feature frequencies are more even (better not to ignore features) and features are sparse in data (error correction is easier).
We can see that in Figure 1, where features are sparse in data, the losses in the strong superposition regime are indeed much smaller than those in the weak superposition regime across several feature frequency distributions.
Figure 3: Weight decay can tune the degree of superposition. (a) Positive weight decay ($\gamma=1$ in the figure) has $\|W_{i}\|_{2}$ near 0 or 1, with frequent features more likely to be represented (color means $\|W_{i}\|_{2}$ in frequency-rank plots). Negative weight decay ($\gamma=-1$) has $\|W_{i}\|_{2}$ around 11. We show results when $\alpha=1,~m=100$, yet the claim is generally true. (b) For all models, small weight decays lead to strong superposition, and large weight decays lead to no superposition ($\phi_{1/2}\approx m/n$). More data in Appendix D.3.
If one regime is more preferred, we want to approach it more quickly in training. If it is not preferred, we also want to study the scaling behaviors scientifically in that regime. To this end, we introduce a decoupled weight decay (or growth) term in training to tune the degree of superposition:
@@
where $\eta_{t}$ is the learning rate and Wi,tW{i,t} is the $i$th row of the weight matrix at step $t$ (vector operations are element-wise). For weight decay $\gamma<0$, the update corresponds to gradient descent on $(|W_{i,t}|_{2}-1)^{2}$, encouraging unit-norm rows. We implement this weight decay in AdamW [28] optimizer with a warm-up and cosine decay learning rate schedule (details in Appendix B). At each training step, we sample new data.
We find that the weight decay can robustly control superposition. We first see that important features tend to be represented (associated $\|W_{i}\|_{2}>0$), and norms of $W_{i}$ become bimodal, clustering near 0 or 1 (Figure 3a). This allows us to define the fraction of represented features as
@@
namely, the fraction of rows with norm larger than 1/21/2.222In theory, we should use 0 as the threshold. The choice, $1/2$, may minimize misclassifications since norms are near 0 or $1$. Our result is robust to this threshold since norms are very concentrated. We found that weight decay can tune superposition for all models we trained, with small weight decay $\gamma$ giving strong superposition, i.e., $\phi_{1/2}\approx 1\gg m/n$, and large weight decay corresponding to weak superposition, i.e., $\phi_{1/2}\sim m/n$ (Figure 3b). The ability of weight decay to tune superposition is robust to feature frequency distributions (Appendix D.3). We can then systematically study scaling behaviors in different regimes.
The toy model differs from LLMs in architecture, data, and loss. Since we focus on representations rather than next-token prediction, we omit transformer layers. Conceptually, LLMs map a document to a token, with inputs and outputs in different spaces, while the toy model operates within a single shared space. Despite this, the toy model captures key aspects of language structure through engineered sparsity and feature importance, making its data structure aligned with that of LLMs at a high level. While LLMs use cross-entropy loss and the toy model uses squared error, we can show that this does not affect the scaling behaviors (Appendix A.2). Thus, the toy model is a suitable abstraction for studying representation-limited scaling.
@@ 3 Results
For a systematic scan, we set $p_{i}\propto 1/i^{\alpha}$ in this section and can vary the data exponent $\alpha$ to change how skewed $p_{i}$ is. [^3] The word or phrase frequency in natural language follows Zipf’s law, which is a power law ($\alpha=1$). The activation density $E$ is set as $1$, whose value can be shown to not affect the scaling (Appendix D.4). We fix data dimension $n=1000$, vary model dimension $m$ from $10$ to $100$, and sweep weight decay $\gamma$ from 1-1 to $1$. We fit final test losses as a power law, $L\propto 1/m^{\alpha_{m}}$, and call $\alpha_{m}$ the model exponent. More details on hyperparameters are in Appendix B.2.
3.1 Weak superposition